OpenStudy (anonymous):

How can we integrate (1+tan^2(x)) with respect to x if the range is 0 to (pi/4)?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}(1+\tan^2(x))\] - Like this

OpenStudy (anonymous):

[tanx + C] [tan pi/4 - tan 0] =1

OpenStudy (anonymous):

helps to know that \[1+\tan^2(x)=\sec^2(x)\] and that \[\frac{d}{dx}\tan(x)=sec^2(x)\] so the anti derivative of \[1+\tan^2(x)\]is \[\tan(x)\]

OpenStudy (anonymous):

1+tan^2 (x)=sec^2 (x)

OpenStudy (anonymous):

Awesome explanation, satellite73. Thanks!!

OpenStudy (anonymous):

And you too, Sarkar!